# BST: Badminton Stroke-type Transformer for Skeleton-based Action Recognition in Racket Sports

### v2

@@ 4 Experiments
In this section, we present the dataset used in our experiments and compare our results with those obtained by several existing methods, including ST-GCN [33], BlockGCN [37], SkateFormer [9], ProtoGCN [22], and TemPose [16].
Model Publication Modality SP PP Acc Macro-F1 Acc-2 NumParams
ST-GCN [33] AAAI 2018 J × × 0.728 0.656 0.897 3.08M
BlockGCN [37] CVPR 2024 J × × 0.715 0.627 0.886 1.50M
SkateFormer [9] ECCV 2024 J × × 0.712 0.627 0.859 2.38M
ProtoGCN [22] CVPR 2025 J × × 0.723 0.637 0.897 4.11M
TemPose-V [16] CVPRW 2023 J × × 0.720 0.642 0.892 1.62M
TemPose-V [16] CVPRW 2023 J+B × × 0.730 0.645 0.895 1.62M
TemPose-PF J+B × 0.738 0.677 0.905 1.68M
TemPose-SF J+B × 0.7602 0.6922 0.9196 1.65M
BST-0 J+B × 0.7626 0.6908 0.9206 1.81M
TemPose-TF [16] CVPRW 2023 J+B 0.7683 0.7027 0.9242 1.71M
BST-1 J+B 0.7681 0.7004 0.9252 1.85M
BST-2 J+B 0.7689 0.7023 0.9260 1.79M
BST-3 J+B 0.7695 0.7043 0.9267 1.85M
Table 1: Model comparison on the testing set without our segmentation strategy. J and B in the modality denote the players’ joints and bones. SP and PP represent additional inputs of shuttlecock and player positions. TemPose-PF and TemPose-SF are intermediate products derived by us from the original model, positioned between TemPose-V and TemPose-TF. The suffix ”F” denotes the fusion of the additional inputs. Acc-2 denotes top-2 accuracy.
Model Modality SP PP Acc Macro-F1 Acc-2
@@
BST-1 J+B 0.7704 0.7026 0.9355
BST-2 J+B 0.7687 0.7000 0.9312
BST-3 J+B 0.7710 0.7042 0.9340
Table 2: Model comparison on the testing set after our segmentation strategy.
Model Modality SP PP Acc Macro-F1 Acc-2
TemPose-SF J+B × 0.591 0.453 0.781
BST-0 J+B × 0.642 0.512 0.812
TemPose-TF [16] J+B 0.601 0.445 0.775
BST-1 J+B 0.640 0.538 0.828
BST-2 J+B 0.636 0.518 0.820
BST-3 J+B 0.640 0.534 0.825
Table 3: Model comparison trained on 25% of the training set without our segmentation strategy.
4.1 Dataset
ShuttleSet333 ShuttleSet has labels and video links. These videos are on BWF TV (https://www.youtube.com/@bwftv).  [32], the largest publicly available badminton video dataset, contains 104 sets, 3,685 rallies, and 36,492 strokes across 44 matches played between 2018 and 2021, featuring 27 top-ranking male and female singles players. After excluding incorrectly labeled and problematic data, a total of 40 matches, 33429 strokes remained for analysis. We set 30 matches for training, 5 matches for validation, and 5 matches for testing.
The original ShuttleSet dataset consists of 19 distinct stroke categories, including a ”none” type. One category, driven flight, containing fewer than 50 strokes across the entire dataset was merged into the ”none” category by us. Additionally, since the dataset does not differentiate between strokes played by the top or bottom player within each category (except for ”none”), the number of stroke categories (except for ”none”) was doubled. This adjustment resulted in a total of 35 categories, including ”none”. (See Tab. A for the complete list of classes.)
4.2 Main Results
Tab. 1 demonstrates that even the most advanced SAR models struggle to achieve satisfactory accuracy. Without incorporating additional inputs, such as shuttlecock trajectories and player positions, significant performance improvements are difficult to attain. However, integrating such information is not straightforward for these models. This highlights the need for specialized models like TemPose [16], which are specifically designed to handle these additional inputs.
Initially, by focusing solely on player position fusion, TemPose-PF achieves results that surpass those of TemPose-V. The reason for this improvement is that player positions imply some information about the shuttlecock trajectory, that is, the players are always chasing the shuttlecock during a rally. On the other hand, integrating only shuttlecock positions, TemPose-SF leads to a substantial improvement of over two percentage points. Furthermore, combining both types of fusion further enhances performance. Notably, our architecture BST, achieves a modest improvement over TemPose when using a fixed-width clipping strategy.
Finally, the incorporation of our segmentation strategy results in even greater performance gains on the metrics across BST models, significantly exceeding the performance of TemPose using our segmentation strategy, as shown in Tab. 2.
For a deeper look into the results between Tab. 1 and Tab. 2, we can observe two phenomena. First, the performance of TemPose drops when using our segmentation strategy. Our segmentation strategy creates variable sequence lengths due to the fluctuating match pace. TemPose struggles with these, possibly due to limitations in its Interactional Transformer, which is designed to process people as a sequence. (We have mentioned this in the last paragraph of Sec. 3.3.1.) The other phenomenon is that when comparing the results of BSTs themselves, the differences between BSTs (without vs. with our segmentation strategy) from BST-1 to BST-3 are [0.23%, 0.22%, 1.03%], [-0.02%, -0.23%, 0.52%], [0.15%, -0.01%, 0.73%]. Notably, 2 of the 3 observed performance drops occur with BST-2. This is because BST-2’s pathways, prior to the MLP head, are limited to two branches (trajectory information from player 1 and from player 2), lacking the overall trajectory information. Furthermore, our segmentation strategy primarily focuses on shuttlecock trajectory information. Consequently, the performance drop is more pronounced for BST-2 than for BST-3. Despite a modest drop on 3 out of 9 metrics, BSTs still benefit from our segmentation strategy, especially with high class granularity (See Sec. 5.1 for more details), where Acc-2 becomes more important. Moreover, compared to the best of TemPose in Tab. 1, BSTs outperform it on 7 out of 9 metrics, with one loss being only -0.01%. Lastly, our prototype BST-0, even without PP input but utilizing our segmentation strategy, still beats the best of TemPose on Acc-2 (0.9306 vs. 0.9242).
4.3 Results on Reduced Training Set
Tab. 3 shows the results of models trained on only (each class) 25% of the training set and tested on the same testing set. This indicates that our BST architecture demonstrates superior generalization capability compared to TemPose [16] when the training dataset is small and reference information is limited.
@@ 4.3 Results on Reduced Training Set
Figure 7: Normalized confusion matrices of BST-0 on our dataset. The sum of the elements in each column in the left matrix is equal to 1, and the sum of the elements in each row in the right matrix is equal to 1.
@@
Figure 8: Example of a 2D pose vs. 3D pose in a badminton video frame. The ninth joint represents the human nose. The red arrows indicate the directions the bottom player was facing. We can see the bottom player was facing her opponent in fact, but the 3D pose exhibits a clear error in the facing direction. We suspect that this 3D model will often assume that if both forearms are pointing roughly forward, it will lead the model to assume that the face is also pointing in that direction. (The visualizations in (b), (c), and (d) are generated by the Mayavi [27] python package.)
Model Accuracy Macro-F1 Accuracy-2
TemPose-SF 0.7602 / 0.7458 0.6922 / 0.6696 0.9196 / 0.9066
BST-0 0.7626 / 0.7498 0.6908 / 0.6768 0.9206 / 0.9080
TemPose-TF [16] 0.7683 / 0.7563 0.7027 / 0.6852 0.9242 / 0.9164
BST-3 0.7695 / 0.7560 0.7043 / 0.6806 0.9267 / 0.9155
@@
Table 4: Model with 2D vs. 3D poses comparison on the testing set without our segmentation strategy. The value on the left side of the slash is the result of the model with 2D poses, and the right side is the result of the model with 3D poses.
5 Discussion
5.1 Classifying Difficulty on ShuttleSet
Although our model achieved the state-of-the-art performance on ShuttleSet [32], the accuracy of it was still not high enough. We suspect that this is primarily due to the excessive granularity of certain categories. The normalized confusion matrices of BST-0 are shown in Fig. 7. The 0th to 16th strokes are executed by the top player, the 17th to 33rd strokes are hit by the bottom player, and the 34th stroke belongs to the ”none” category. Observing on both confusion matrices, the model struggles to differentiate between similar stroke-types, such as the 2nd and 3rd strokes (top smash and top wrist smash) and the 19th and 20th strokes (bottom smash and bottom wrist smash). A closer inspection reveals that the model often misclassifies the 1st stroke (top return net) as the 13th stroke (top defensive return drive) that shown in the left matrix. Additionally viewing the diagonal elements in the right matrix, the 13th stroke (top defensive return drive) has the worst recall. These results suggest that the model has difficulty distinguishing between subtle variations in the strokes with similar hitting characteristics. We observed the same phenomenon when employing other models as well.
5.2 Player 2D Joints vs. 3D Joints
@@ 5.1 Classifying Difficulty on ShuttleSet
While we explored using 3D player joints as input to our model, the performance was inferior to that achieved with 2D joints. We attribute this to the fact that the HPE model is trained on a general HPE dataset, encompassing a broader range of poses than those specific to badminton players. Additionally, the 3D joints are estimated from the 2D joints, which may further contribute to a bias towards general human poses, rather than the specialized poses of badminton players. We can see an example of this in Fig. 8. This issue would mislead the model to predict the wrong action, since the 3D pose is not accurate enough to represent the player’s pose in the badminton game, as shown in Tab. 4.
5.3 Limitations and Future Work
@@
Our method has an obvious limitation: it relies on accurate hit frame detection and shuttlecock trajectory tracking models. Fortunately, these issues are expected to be resolved in the future, similar to how the transition from HAR to SAR was facilitated by advanced HPE models, which enabled SAR to perform more effectively. Moreover, the existing 2D shuttlecock trajectory tracking models are quite mature, and we believe that there will be more 3D tracking models in the future. Although this paper primarily focuses on stroke classification in singles badminton, the potential applicability of our method to other racket sports, such as tennis and table tennis, or even doubles, is worth exploring.
6 Conclusion
In this paper, we first introduce a novel segmentation strategy that segments a match video into clips containing two to three badminton strokes. We propose BST for badminton stroke classification. Through experiments, we show that our method outperforms existing state-of-the-art methods, highlighting the importance of the shuttlecock information.
Supplementary Material
A Training Speed
On a NVIDIA RTX 4070 Ti Super setup, training BST-3 takes 29 sec per epoch, while training TemPose-TF takes 31 sec per epoch. Fig. A also shows that training BST-3 converges faster than TemPose-TF.
Figure A: Loss Curves of BST-3 and TemPose-TF.
B Training Settings

### v3

@@ 4 Experiments
In this section, we present the three datasets used in our experiments and compare our results with those obtained by several existing methods, including ST-GCN [35], BlockGCN [40], SkateFormer [10], ProtoGCN [24], and TemPose [18].
Model Modality SP $P$ Acc Macro-F1 Acc-2
TemPose-SF J+B × 0.6504 0.5518 0.8262
BST-0 J+B × 0.6946 0.6248 0.8566
TemPose-TF [18] J+B 0.6636 0.5904 0.8344
BST J+B 0.6884 0.6196 0.8552
BST-CG J+B 0.7020 0.6334 0.8672
BST-AP J+B 0.7038 0.6302 0.8584
BST-CG-AP J+B 0.6924 0.6238 0.8566
Table 3: Model comparison trained on 25% ShuttleSet (25 classes) with fixed-width strategy.
Model Publication Modality SP $P$ Acc Macro-F1 Acc-2
ST-GCN [35] AAAI 2018 J × × 0.5602 0.4870 0.7732
BlockGCN [40] CVPR 2024 J × × 0.5782 0.4958 0.7970
ProtoGCN [24] CVPR 2025 J × × 0.5656 0.4798 0.7888
TemPose-V [18] CVPRW 2023 J × × 0.5708 0.5248 0.8134
TemPose-V [18] CVPRW 2023 J+B × × 0.5764 0.5138 0.8308
TemPose-PF J+B × 0.5986 0.5516 0.8536
TemPose-SF J+B × 0.6204 0.5440 0.8380
BST-0 J+B × 0.6460 0.5671 0.8569
TemPose-TF [18] CVPRW 2023 J+B 0.6199 0.5448 0.8517
BST J+B 0.6420 0.5600 0.8620
BST-CG J+B 0.6437 0.5734 0.8639
BST-AP J+B 0.6517 0.5799 0.8628
BST-CG-AP J+B 0.6392 0.5580 0.8574
Table 4: Model comparison on BadmintonDB (18 classes) with predefined strategy.
Model Publication Modality SP $P$ Acc Macro-F1 Min-F1 Acc-2
ST-GCN [35] AAAI 2018 J × × 0.9890 0.9882 0.9766 0.9956
TemPose-V [18] CVPRW 2023 J × × 0.9824 0.9822 0.9732 0.9946
TemPose-V [18] CVPRW 2023 J+B × × 0.9780 0.9770 0.9622 0.9922
TemPose-PF J+B × 0.9844 0.9846 0.9688 0.9948
TemPose-SF J+B × 0.9766 0.9765 0.9608 0.9938
BST-0 J+B × 0.9904 0.9901 0.9807 0.9992
TemPose-TF [18] CVPRW 2023 J+B 0.9866 0.9872 0.9767 0.9956
BST J+B 0.9890 0.9889 0.9795 0.9978
BST-CG J+B 0.9913 0.9913 0.9829 0.9986
BST-AP J+B 0.9923 0.9922 0.9845 0.9990
BST-CG-AP J+B 0.9878 0.9879 0.9783 0.9970
Table 5: Model comparison on TenniSet (6 classes) with predefined strategy.
4.1 Datasets
ShuttleSet [^2] ShuttleSet has labels and video links. These videos are on BWF TV (https://www.youtube.com/@bwftv).  [34], the largest publicly available badminton video dataset, contains 104 sets, 3,685 rallies, and 36,492 strokes across 44 matches played between 2018 and 2021, featuring 27 top-ranking male and female singles players. After excluding incorrectly labeled and problematic data, a total of 40 matches, 33481 strokes remained for analysis. We set 30 matches for training, 5 matches for validation, and 5 matches for testing.
BadmintonDB [3] contains 7658 strokes across 8 badminton singles matches after data cleaning. These matches were played by two players, Kento Momota and Anthony Sinisuka Ginting, in 2018 and 2019. The time period of each clip is predefined, and on our observation, the target hit frames are always near the end of the clips (about half of them even after the end). This means that, usually, the information in stage 1 of our clipping strategy is utilized only. We split the dataset into 8:1:1 for training, validation, and testing on each class.
TenniSet [12] contains 3569 strokes across 5 tennis singles matches. The time period of each clip is also predefined. We followed the split configuration predefined in this dataset for training, validation, and testing.
4.2 Results
Results on ShuttleSet (merged). [^3] Several stroke-types are merged. Refer to the supplementary material for details. Tab. 1 demonstrates that even the most advanced SAR models struggle to achieve satisfactory accuracy. Without incorporating additional inputs, such as shuttlecock trajectories and player positions, significant performance improvements are difficult to attain. However, integrating such information is not straightforward for these models. This highlights the need for specialized models like TemPose [18], which are specifically designed to handle these additional inputs.
Initially, by focusing solely on player position fusion, TemPose-PF achieves results that surpass those of TemPose-V. The reason for this improvement is that player positions imply some information about the shuttlecock trajectory, that is, the players are always chasing the shuttlecock during a rally. On the other hand, integrating only shuttlecock positions, TemPose-SF achieves a greater improvement over TemPose-V than TemPose-PF does. Furthermore, combining both types of fusion further enhances performance. Notably, our BST series achieves some improvement over TemPose on 18/20 metrics when using the fixed-width clipping strategy.
Moreover, our clipping strategy yields even greater performance gains for all TemPose and BST series models, as shown in Tab. 2.
Finally, we also conducted experiments using only (each class) 25% of the training set and tested on the same testing set, as shown in Tab. 3. The results indicate that our BST architecture demonstrates superior generalization capability compared to TemPose when the training set is small and reference information is limited.
Results on BadmintonDB. As demonstrated in Tab. 4, these advanced SAR models still struggle to achieve satisfactory accuracy. While the TemPose-SF outperforms TemPose-PF on accuracy, it falls short on macro-F1 and top-2 accuracy. Furthermore, we found that building upon an SP with the addition of PP resulted in performance drops on several metrics (e.g. TemPose-SF vs. TemPose-TF and BST-0 vs. BST). Nevertheless, our BST series consistently outperforms previous methods.
Results on TenniSet. This dataset is less challenging than the previous two, primarily due to its reduced number of categories with a more balanced distribution. As shown in Tab. 5, all models achieve very high accuracy, macro-F1, top-2 accuracy, and even min-F1. While adding bones information slightly yields a modest improvement on the previous two datasets, it results in a slight performance drop on this dataset. Despite this, we still used the J+B modality on the subsequent models for consistency. This time, ST-GCN outperforms TemPose models on most metrics. However, our BST series still achieves the best results across most metrics.
5 Conclusion
In this paper, we first introduce a novel clipping strategy that segments a match video into clips containing highly relevant frames for the target stroke. We propose BST architecture, including its backbone BST-0 and several enhancing modules, for badminton stroke classification. Through experiments, we show that our method outperforms existing state-of-the-art methods not only in two badminton datasets but also in a tennis dataset, highlighting the importance of the ball information. A detailed comparison of model training speed is provided in the supplementary material, along with a comprehensive analysis of performance using 2D vs. 3D joints.
Supplementary Material
Contents
Appendix A Training Speed
On a NVIDIA RTX 4090 setup, BST-CG-AP trains slightly faster than TemPose-TF, as shown in Tab. A.
Model Seq Len Ep. Speed #Param
ST-GCN [35] 30 11.1 s 3.08M
BlockGCN [40] 30 111.6 s 1.50M
SkateFormer [10] 30 41.8 s 2.38M
ProtoGCN [24] 30 42.4 s 4.11M
TemPose-TF* [18] 100 11.0 s $\sim$1.71M
TemPose-TF [18] 100 10.0 s 1.71M
BST-CG-AP 100 9.7 s 1.85M
Table A: Model training speed per epoch on ShuttleSet and the numbers of their parameters. TemPose-TF* is the original implementation version.
Fig. A also shows that training BST-CG-AP converges faster than TemPose-TF. The performance gap between TemPose-TF* and TemPose-TF is likely due to the authors’ use of a single TCN object to handle both player positions and the shuttlecock.
Figure A: Loss Curves of BST-CG-AP, TemPose-TF and TemPose-TF*.
Appendix B Discussion
Model Publication Modality SP $P$ Acc Macro-F1 Acc-2
ST-GCN [35] AAAI 2018 J × × 0.7198 0.6396 0.8890
BlockGCN [40] CVPR 2024 J × × 0.7108 0.6310 0.8822
SkateFormer [10] ECCV 2024 J × × 0.7156 0.6304 0.8588
ProtoGCN [24] CVPR 2025 J × × 0.7152 0.6282 0.8808
TemPose-V [18] CVPRW 2023 J × × 0.7154 0.6326 0.8920
TemPose-V [18] CVPRW 2023 J+B × × 0.7216 0.6378 0.8952
TemPose-PF J+B × 0.7384 0.6670 0.9090
TemPose-SF J+B × 0.7602 0.6922 0.9196
BST-0 J+B × 0.7626 0.6908 0.9206
TemPose-TF [18] CVPRW 2023 J+B 0.7683 0.7027 0.9242
BST J+B 0.7676 0.6966 0.9236
BST-CG J+B 0.7681 0.7004 0.9252
BST-AP J+B 0.7689 0.7023 0.9260
BST-CG-AP J+B 0.7695 0.7043 0.9267
Table B: Model comparison on ShuttleSet (35 classes) with fixed-width strategy.
Model Modality SP $P$ Acc Macro-F1 Acc-2
@@ Appendix B Discussion
BST J+B 0.7698 0.7010 0.9331
BST-CG J+B 0.7704 0.7026 0.9355
BST-AP J+B 0.7687 0.7000 0.9312
BST-CG-AP J+B 0.7710 0.7042 0.9340
Table C: Model comparison on ShuttleSet (35 classes) with our clipping strategy.
@@
Figure B: Normalized confusion matrices of BST-0 on ShuttleSet (35 classes). The sum of the elements in each column in the left matrix is equal to 1, and the sum of the elements in each row in the right matrix is equal to 1.
Model Modality SP $P$ Acc Macro-F1 Acc-2 TemPose-SF J+B × 0.5888 0.4552 0.7760 BST-0 J+B × 0.6264 0.5136 0.8098 TemPose-TF [18] J+B 0.5822 0.4372 0.7560 BST J+B 0.6323 0.5238 0.8202 BST-CG J+B 0.6436 0.5266 0.8234 BST-AP J+B 0.6374 0.5276 0.8164 BST-CG-AP J+B 0.6338 0.5204 0.8218
Table D: Model comparison trained on 25% ShuttleSet (35 classes) with fixed-width strategy.
B.1 Classifying Difficulty on ShuttleSet
At the beginning of this work, we ran the experiments on ShuttleSet [34] (35 classes), but the accuracy of it was not high enough, as shown in Tab. B and Tab. C. We suspect that this is primarily due to the excessive granularity of certain categories. The normalized confusion matrices of BST-0 are shown in Fig. B. The 0th to 16th strokes are executed by the top player, the 17th to 33rd strokes are hit by the bottom player, and the 34th stroke belongs to the ”none” category. Observing on both confusion matrices, the model struggles to differentiate between similar stroke-types, such as the 2nd and 3rd strokes (top smash and top wrist smash) and the 19th and 20th strokes (bottom smash and bottom wrist smash). A closer inspection reveals that the model often misclassifies the 1st stroke (top return net) as the 13th stroke (top defensive return drive) that is shown in the left matrix. Additionally viewing the diagonal elements in the right matrix, the 13th stroke (top defensive return drive) has the worst recall. These results suggest that the model has difficulty distinguishing between subtle variations in the strokes with similar hitting characteristics. We observed the same phenomenon when employing other models as well.
For a deeper look into the results between Tab. B and Tab. C, we can observe two phenomena. First, the performance of TemPose drops when using our clipping strategy. Our clipping strategy creates variable sequence lengths due to the fluctuating match pace. TemPose struggles with these, possibly due to limitations in its Interactional Transformer, which is designed to process people as a sequence. (We have mentioned this in the last paragraph of Sec. 3.3.1.) The other phenomenon is that when comparing the results of BSTs themselves, the differences between BSTs (fixed-width vs. our clipping strategy) from BST-CG to BST-CG-AP are [0.23%, 0.22%, 1.03%], [-0.02%, -0.23%, 0.52%], [0.15%, -0.01%, 0.73%]. Notably, 2 of the 3 observed performance drops occur with BST-AP. This is probably because BST-AP’s pathways, prior to the MLP head, are limited to two branches (information from the blue player and from the green player), lacking the overall trajectory information. Furthermore, our clipping strategy primarily focuses on shuttlecock trajectory information. Consequently, the performance drop is more pronounced for BST-AP than for BST-CG-AP. Despite a modest drop on 3 out of 9 metrics, BSTs still benefit from our clipping strategy, especially with high class granularity, where Acc-2 becomes more important. Moreover, compared to the best of TemPose in Tab. B, BSTs outperform it on 7 out of 9 metrics, with one loss being only -0.01%. Lastly, our backbone BST-0, even without PP input but utilizing our clipping strategy, still beats the best of TemPose on Acc-2 (0.9306 vs. 0.9242).
Tab. D shows the results of models trained on only (each class) 25% of the training set and tested on the same testing set. This indicates that our BST architecture still demonstrates superior generalization capability compared to TemPose when the training dataset is small and reference information is limited, even with excessive class granularity.
B.2 Player 2D Joints vs. 3D Joints
@@ B.1 Classifying Difficulty on ShuttleSet
Figure C: Example of a 2D pose vs. 3D pose in a badminton video frame. The ninth joint represents the human nose. The red arrows indicate the directions the bottom player was facing. We can see the bottom player was facing her opponent in fact, but the 3D pose exhibits a clear error in the facing direction. We suspect that this 3D model will often assume that if both forearms are pointing roughly forward, it will lead the model to assume that the face is also pointing in that direction. (The visualizations in (b), (c), and (d) are generated by the Mayavi [29] python package.)
Model Accuracy Macro-F1 Accuracy-2 TemPose-SF 0.7602 / 0.7432 0.6922 / 0.6670 0.9196 / 0.9062 BST-0 0.7626 / 0.7456 0.6908 / 0.6696 0.9206 / 0.9102 TemPose-TF [18] 0.7683 / 0.7579 0.7027 / 0.6866 0.9242 / 0.9176 BST-CG-AP 0.7695 / 0.7525 0.7043 / 0.6760 0.9267 / 0.9181
@@
Table E: Model with 2D vs. 3D poses comparison on ShuttleSet (35 classes) with fixed-width clipping strategy. The value on the left side of the slash is the result of the model with 2D poses, and the right side is the result of the model with 3D poses.
@@
While we explored using 3D player joints as input to our model, the performance was inferior to that achieved with 2D joints. We attribute this to the fact that the HPE model is trained on a general HPE dataset, encompassing a broader range of poses than those specific to badminton players. Additionally, the 3D joints are estimated from the 2D joints, which may further contribute to a bias towards general human poses, rather than the specialized poses of badminton players. We can see an example of this in Fig. C. This issue would mislead the model to predict the wrong action, since the 3D pose is not accurate enough to represent the player’s pose in the badminton game, as shown in Tab. E.
B.3 Limitations and Future Work
@@
Our method has an obvious limitation: it relies on accurate hit frame detection and shuttlecock trajectory tracking models. Fortunately, these issues are expected to be resolved in the future, similar to how the transition from HAR to SAR was facilitated by advanced HPE models, which enabled SAR to perform more effectively. Moreover, the existing 2D shuttlecock trajectory tracking models are quite mature, and we believe that there will be more 3D tracking models in the future.
Although this paper primarily focuses on stroke classification in singles badminton, and also experiments on tennis, the potential applicability of our method to other racket sports, such as table tennis, or even doubles, is worth exploring.
Appendix C Details in Clipping
C.1 Source of the Hit Frame and Rally Information
Since the clipping strategies are predefined in BadmintonDB [3] and TenniSet [12], and the hit frame and rally information are provided in ShuttleSet [34], we don’t need to detect hit frames and rallies by ourselves. If these were not provided, we could use the HitNet model proposed in MonoTrack [25] to detect hit frames. To detect rallies, we note that time intervals between successive rallies are in general much larger than the time differences between two typical successive hit frames in one rally. For example, the time difference between the last hit frame number of a rally and the first hit frame number of the next rally must be larger than that between the two consecutive hit frames in the same rally, which can be presented as $h^{(i+1)}_{1}-h^{(i)}_{n^{(i)}}>h^{(k)}_{j+1}-h^{(k)}_{j},\forall\ 1\leq k\leq r$, where $r$ is the total number of rallies in the match video.
C.2 Practical Settings
In practice, we set some limits on $\hat{h}^{(i)}_{j-1}$ and $\hat{h}^{(i)}_{j+1}$ in our strategy to prevent the model from accessing irrelevant information caused by remaining unclean data. Specifically, the actual time distance of these two frame numbers from the target frame number does not exceed 1.5 seconds (before adding $\epsilon$). Additionally, $t$ is set to the half of the FPS of the video representing 0.5 seconds interval, and $\epsilon$ is set to $\frac{t}{2}$.
Appendix D Training Settings